# Comparing Two Proxy Methods for Causal Identification

### v2

@@ 3 Causal Target Identification With Proxies via Array Decomposition
Another prominent strategy for identifying counterfactual quantities with proxies of hidden variables employs array decomposition to recover the full law of hidden and observed variables, up to the labels of hidden variable categories (e.g., Kuroki and Pearl (2014); Allman et al. (2015); Zhou and Tchetgen (2024)). We now review the assumptions of this approach in the context of the discrete hidden confounding model depicted in Figure 3, from Kuroki and Pearl (2014). In Appendix C, we provide analogous assumptions which leverage proxies of a hidden mediator in the absence of hidden confounder proxies, enabling identification causal effects via the frontdoor adjustment formula in Equation 7.
In the model depicted in Figure 3, taken to be fully discrete for now, where $Z$ and $W$ function as proxy variables for the unobserved confounder $U$, Kuroki and Pearl (2014) show that the latent components $f_{Y\mid U,A}(y\mid u_{i},a)$ and $f_{U}(u_{i})$ are identifiable jointly up to label of the latent categories of $U$. With these components identified, Kuroki and Pearl (2014) recover $f_{Y(a)}(y)$ from the adjustment formula in Equation 1. We focus on the fully discrete version of the model for the time being, deferring discussion of continuous models to a later point.
@@
Kuroki and Pearl (2014) employ the following assumptions:
Define for each $a\in\mathcal{A}$,
$\mathbf{P}_{Z\mid U,a}=\big[\,f_{Z\mid U,A}(z\mid u_{1},a),\;\dots,\;f_{Z\mid U,A}(z\mid u_{|\mathcal{U}|},a)\,\big],$
$\mathbf{P}_{W\mid U}=\big[\,f_{W\mid U}(w\mid u_{1}),\;\dots,\;f_{W\mid U}(w\mid u_{|\mathcal{U}|})\,\big],$
$\mathbf{P}_{Y\mid U,a}=\big[\,f_{Y\mid U,A}(y\mid u_{1},a),\;\dots,\;f_{Y\mid U,A}(y\mid u_{|\mathcal{U}|},a)\,\big].$
@@
$W,Z,Y,\text{ are mutually independent }\mid U,A.$
@@
For each $a\in\mathcal{A}$, some row of $\mathbf{P}_{Y\mid U,a}$ has distinct entries.
A detailed proof of identification is given in Kuroki and Pearl (2014). We provide a short summary in Appendix B. For Assumption 10 to be satisfied, $Z$ and $W$ must have the same number of levels. Assumption 10 can be relaxed to left invertibility of $\mathbf{P}_{Z\mid U,a}$ and $\mathbf{P}_{W\mid U}$, allowing cases where $Z$ and $W$ each have at least as many levels as $U$. Perhaps surprisingly, Assumption 10 can even be relaxed to allow settings where both proxies $Z$ and $W$ have fewer levels than the latent variable. These generalizations are discussed in Section 3.1.
Techniques for identifying causal targets via array decomposition, presented in works including Kuroki and Pearl (2014), Allman et al. (2015), and Zhou and Tchetgen (2024) can be viewed as special cases of Kruskal’s uniqueness theorem for the Candecomp/Parafac (CP) decomposition (Kruskal, 1977), which we discuss in Section 3.1. As a consequence, any factor of the full law of hidden and observed variables can be recovered in these settings up to latent category labels. A natural extension of these ideas to continuous settings is explored in Section 3.2.
@@ 3.1 Discrete Hidden Variable Models
The array decomposition approach in discrete models is a special case of Kruskal’s uniqueness theorem for the Candecomp/Parafac (CP) decomposition of a three-way array (or data tensor, informally), as noted in Allman et al. (2015). The theorem provides conditions under which any factor of the full law of hidden and observed variables is identified up to latent state labels, and rules out any lower-dimensional representation of the latent variable (Kruskal, 1977).
The remainder of this section presents the theorem and examines in detail how its assumptions relate to those of Kuroki and Pearl (2014). A detailed and concise proof of the theorem is provided in Rhodes (2010).
@@
The CP decomposition of a three-way array $\boldsymbol{\mathcal{T}}$ with dimensions $M\times N\times J$ consists of three factor matrices $\mathbf{A}\in\mathbb{R}^{M\times R}$, $\mathbf{B}\in\mathbb{R}^{N\times R}$, and $\mathbf{C}\in\mathbb{R}^{J\times R}$ such that the $j^{\text{th}}$ slice of $\boldsymbol{\mathcal{T}}$ can be written as (Stegeman and Sidiropoulos, 2007; Kruskal, 1977):
$\boldsymbol{\mathcal{T}}_{j}=\mathbf{A}\mathbf{C}_{j}\mathbf{B}^{T},\text{where }\mathbf{C}_{j}\text{ is diagonal}$ (5)
[…]
@@
The three-way rank of $\boldsymbol{\mathcal{T}}$ is the smallest integer $R$ for which its CP decomposition in equation (5) holds.
@@
The Kruskal rank, or $k$-rank, of a matrix is the largest integer $k$ such that any subset of $k$ columns is linearly independent.
@@
A three-way array $\boldsymbol{\mathcal{T}}$ is essentially unique and has three-way rank R if the following condition holds:
@@
By Theorem 3.5, identification of latent factors
$f_{W\mid U,A}(w\mid u_{i},a),f_{Y\mid U,A}(y\mid u_{i},a),\text{ and }f_{U,Z\mid A}(u_{i},z\mid a)$
reduces to solving the decomposition for each $a\in\mathcal{A}$. Further by $W\perp U\mid A$, latent factors for label $i$ recovered across values of $a\in\mathcal{A}$ can be aligned. Hence, the full law $f_{A,Y,U,W,Z}(a,y,u_{i},w,z)$ is identified up to labels $i$.
The original proof of Theorem 3.5 is nonconstructive, and can be viewed in terms of uniquely determined eigenspaces (Rhodes, 2010). In the special setting considered by Kuroki and Pearl (2014), the authors showed latent components can be obtained directly via eigendecomposition tasks. Indeed, when $k_{A}=k_{B}=R=|\mathcal{U}|$ and $k_{C}\geq 2$, latent components can be constructed from applying the argument from Kuroki and Pearl (2014) to submatrices of $\boldsymbol{\mathcal{T}}_{j}$, as also noted by Allman et al. (2015). Modern methods for recovering latent factors under the general assumptions of Theorem 3.5 often rely on iterative procedures such as alternating least squares or gradient-based optimization (Uschmajew, 2012; Kolda and Hong, 2020).
Notably, Theorem 3.5 leverages constraints across all slices of the three-way array $\boldsymbol{\mathcal{T}}$. Since the inequality treats $W$, $Z$, and $Y$ symmetrically, variation in the conditional laws $f_{Y\mid U,A}(y\mid u_{i},a)$ across levels of $U$ can offset limited variation in laws $f_{W\mid U}(w\mid u_{i})$ or $f_{Z\mid U,A}(z\mid u_{i},a)$. In this manner, the inequality in Theorem 3.5 allows for situations where $Z$ and/or $W$ have fewer levels than $U$, as long as $Y$ has enough levels. A necessary requirement for the inequality in Theorem 3.5 to hold is that variables $W$, $Z$, and $Y$ collectively have at least $2|\mathcal{U}|+2$ distinct categories.
Note the CP decomposition framework is typically introduced in the simpler case of a single latent variable with three proxy variables that are mutually independent conditional upon the latent, depicted in Figure 4. The resulting three-way array formed by the proxy variables uniquely determines the full law up to latent labels, given the condition in Theorem 3.5 holds (Kruskal, 1977). We now turn to how this framework generalizes for continuous models.
@@ 3.2 Continuous Hidden Variable Models
For continuous hidden variable models, Hu and Schennach (2008) provide a natural extension of the array decomposition framework for unique determination of the full law up to latent state labels. A key requirement is Assumption 12, which states that three observed variables are mutually independent conditional on the latent variable $L$, as illustrated in Figure 4. Hu and Schennach (2008) also invoke the bounded density conditions in Assumption 13.
@@
$W,Z,Y\text{ are mutually independent }\mid L.$
@@
$W,Z,Y,$ and $L$ admit bounded, nonzero densities with respect to the Lebesgue measure on $\mathcal{W}\times\mathcal{Z}\times\mathcal{L}$, and with respect to some dominating measure $\mu$ on $\mathcal{Y}$. All marginal and conditional densities are assumed to be bounded as well.
LYWZ
Figure 4: Latent Variable with Proxies;
Supports $\mathcal{Y,W,Z,L}$
Additionally, the following two assumptions are imposed, paralleling Assumptions 10 and 11 in Kuroki and Pearl (2014), or the inequality required by Theorem 3.5 from Kruskal (1977).
@@
For $g$ square-integrable, $\mathbb{E}[g(l)\mid w]=0\,\text{ a.s.}\iff g(l)=0\,\text{ a.s.}$ and $\mathbb{E}[g(l)\mid z]=0\,\text{ a.s.}\iff g(l)=0\,\text{ a.s.}$
@@
For all $i\neq j$, the conditional distributions $f_{Y\mid L}(y\mid l_{i})$ and $f_{Y\mid L}(y\mid l_{j})$ differ with positive probability under the marginal distribution of $Y$.
Hu and Schennach (2008) prove that the full law $f_{Y,W,Z,U}(y,w,z,l)$ is uniquely determined up to labeling of latent states [Assumptions 19 or 20 give the labelings]. For ease of comparison, we restate Assumptions 1315 in a form that ensures unique identification of the counterfactual distribution $f_{Y(a)}(y)$ in the setting depicted in Figure 3 from Kuroki and Pearl (2014), instead allowing variables to be continuous. This was proven by Deaner (2023).
@@
$W,Z,Y,U\mid A=a$ have bounded, nonzero densities with respect to the Lebesgue measure on $\mathcal{W}\times\mathcal{Z}\times\mathcal{U}$, and with respect to some dominating measure $\mu$ on $\mathcal{Y}$. All marginal and conditional factors of $f_{W,Z,Y,U\mid a}(w,z,y,u\mid a)$ are bounded as well.
Assumption 17.
For each $a\in\mathcal{A}$ and $g$ square-integrable, $\mathbb{E}[g(u)\mid w,a]=0\,\text{ a.s.}\iff g(u)=0\,\text{ a.s.}$ and $\mathbb{E}[g(u)\mid z,a]=0\,\text{ a.s.}\iff g(u)=0\,\text{ a.s.},$
Assumption 18.
For each $a\in\mathcal{A}$ and $i\neq j$, the conditional distributions $f_{Y\mid U,A}(y\mid u_{i},a)$ and $f_{Y\mid U,A}(y\mid u_{j},a)$ differ with positive probability under the marginal distribution of $Y\mid A=a$.
Assumption 17 is the continuous analogue of Assumption 10 from Kuroki and Pearl (2014), or more generally $k_{A}=k_{B}=R=|\mathcal{U}|$, while Assumption 18 corresponds to Assumption 11 from Kuroki and Pearl (2014), or more generally $k_{C}\geq 2$. Since Theorem 3.5 permits situations where both $Z$ and $W$ have fewer categories than the latent variable $U$ in discrete models, whether the result from Hu and Schennach (2008) admits a similar relaxation remains an open question.
3.3 Identifying Latent Labels
Outside of the hidden confounding context, it is often necessary to impose additional conditions in order to recover the labels of the latent variable. When the treatment is unobserved but proxies of it are available, label recovery for $A$ is essential for identifying a specific counterfactual distribution $f_{Y(a_{i})}(y)=\int f_{Y\mid A,C}(y\mid a_{i},c)\,f_{C}(c)\,dc$ (Zhou and Tchetgen, 2024).
3.3.1 Unbiasedness Conditions for Label Recovery
Hu and Schennach (2008) show that in the model depicted in Figure 4 under Assumptions 1214, and 17, latent state labels can be recovered under an unbiasedness condition, stating that a measure of central tendency of a proxy, conditional on each latent value, must equal the latent value itself. This condition is relayed in Assumption 19.
Assumption 19.
$L$ is ordinal and there exists a known functional $M$ such that for some $V\in\{W,Z,Y\}$ and for each $l\in\mathcal{L},$ $M[f_{V\mid L}(v\mid l)]=l.$
3.3.2 Monotonicity Conditions for Label Recovery
As an alternative to Assumption 19, labels may also be recovered under a monotonicity condition relayed in Assumption 20, stating that a measure of central tendency of a proxy is strictly ordered across latent categories.
Assumption 20.
$L$ is ordinal and there exists a known functional $M$ such that, for some $V\in\{W,Z,Y\}$ and all $l_{i}<l_{j}$, either $M[f_{V\mid L}(v\mid l_{i})]<M[f_{V\mid L}(v\mid l_{j})]$ or $M[f_{V\mid L}(v\mid l_{i})]>M[f_{V\mid L}(v\mid l_{j})].$
@@ 4 Model Comparison
The array decomposition approach in Section 3 can be used to recover counterfactual quantities across a broad class of hidden variable models. For example, Zhou and Tchetgen (2024) apply the method to hidden treatment models. This flexibility underscores a key strength of the array decomposition framework. Moreover, the two proximal strategies for hidden confounding specifically invoke distinct Markov restrictions, which are non-nested, allowing the eigendecompositon approach to swap the roles of $Y$ and AA  (Deaner, 2023).
We now compare the two approaches under a unified set of Markov restrictions compatible with both frameworks. Specifically, we consider the hidden confounding model that is Markov to Figure 3, i.e., reflecting the conditional independences in Assumptions 8 and 9.
Table 1: Comparison for discrete hidden confounding model Markov to Figure 3.
@@
$\mathbf{P}_{Z\mid U,a}$ is left invertible $\exists\mathbf{H}_{a}$ s.t. $\mathbf{P}_{Y\mid Z,a}=\mathbf{H}_{a}\,\mathbf{P}_{W\mid Z,a}$ $\mathbf{P}_{Z\mid U,a}^{-1}$ and $\mathbf{P}_{W\mid U}^{-1}$ exist Some row of $\mathbf{P}_{Y\mid U,a}$ has distinct entries $\mathbf{P}_{Z\mid U,a}^{-1}$ and $\mathbf{P}_{W\mid U}^{-1}$ exist Some row of $\mathbf{P}_{Y\mid U,a}$ has distinct entries
@@
$\mathbf{P}_{Z\mid U,a}$ is left invertible $\exists\mathbf{H}_{a}$ s.t. $\mathbf{P}_{Y\mid Z,a}=\mathbf{H}_{a}\,\mathbf{P}_{W\mid Z,a}$ $2|\mathcal{U}|+2\leq k_{A}+k_{B}+k_{C}$ $\mathbf{P}_{Z\mid U,a}$ is left invertible
@@
$\exists\mathbf{H}_{a}$ s.t. $\mathbf{P}_{Y\mid Z,a}=\mathbf{H}_{a}\,\mathbf{P}_{W\mid Z,a}$ $|\mathcal{U}|+2\leq k_{A}+k_{C}$
@@
Bridge equation approach assumptions are from Miao, Geng and Tchetgen Tchetgen (2018).
@@
4.0.1 Discrete Model Comparison
For discrete models, we compare Assumption 6 and 7 from the bridge equation approach in Miao et al. (2018) with two alternative assumption sets (Table 1). Set 1 is from Kuroki and Pearl (2014); Set 2 is from Kruskal (1977), the generalization of Kuroki and Pearl (2014) via Theorem 3.5.
Comparing Set 1, all model overlap occurs under the assumptions of Kuroki and Pearl (2014). In particular, the invertibility conditions in Assumption 10 from Kuroki and Pearl (2014)—namely, that $\mathbf{P}_{Z\mid U,a}^{-1}$ and $\mathbf{P}_{W\mid U,a}^{-1}$ exist—also imply the existence of $\mathbf{P}_{W\mid Z,a}^{-1}$. This directly satisfies Assumptions 6 and 7 in Miao, Geng and Tchetgen Tchetgen (2018), since one can define
$\mathbf{H}_{a}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{-1}\quad\text{so that}\quad\mathbf{P}_{Y\mid Z,a}=\mathbf{H}_{a}\,\mathbf{P}_{W\mid Z,a}.$
Thus, the model from Kuroki and Pearl (2014) is contained within that of Miao, Geng and Tchetgen Tchetgen (2018).
To compare Set 2, again let
$\mathbf{A}=\mathbf{P}_{W\mid U},\quad\mathbf{B}=\mathbf{P}_{Z,U\mid a},\quad\text{and}\quad\mathbf{C}=\mathbf{P}_{Y\mid U,a},$

### v3

@@ 3 Causal Target Identification With Proxies via Array Decomposition
Another prominent strategy for identifying counterfactual quantities with proxies of hidden variables employs array decomposition to recover the full law of hidden and observed variables, up to the labels of hidden variable values (e.g., Kuroki and Pearl (2014); Allman et al. (2015); Zhou and Tchetgen Tchetgen (2024)). We begin by reviewing this framework in the fully discrete hidden confounding model of Kuroki and Pearl (2014), consistent with the conditional independences depicted in Figure 3. In Appendix D, we provide analogous assumptions for recovering causal effects in a discrete model with proxies of a hidden mediator. We defer discussion of continuous extensions to Section 3.2.
Although the graph in Figure 3 does not impose discreteness, the array decomposition result from Kuroki and Pearl (2014) is most transparently understood in a fully discrete version of this model. In this setting, with $Z$ and $W$ serving as proxies of the unobserved confounder $U$, Kuroki and Pearl (2014) show that the latent components $f_{Y\mid U,A}(y\mid u_{i},a)$ and $f_{U}(u_{i})$ are jointly identifiable up to a permutation of the latent categories of $U$. Once these components are recovered, the counterfactual distribution $f_{Y(a)}(y)$ may be derived from Equation 1.
@@
Kuroki and Pearl (2014) use the following assumptions, adopting the matrix notation $\mathbf{P}_{V\mid T}$ for encoding conditional distributions (see Appendix A), which we use throughout the rest of the paper.
@@
$W,Z,Y\text{ are mutually independent conditional on }U,A.$
@@
For $a\in\mathcal{A}$ and $y\in\mathcal{Y}$, $\Pr(Y=y\mid U=u_{k},A=a)\neq\Pr(Y=y\mid U=u_{j},A=a),\quad k\neq j.$
We provide a short summary of the identification proof given in Kuroki and Pearl (2014) in Appendix C. Assumption 10 can be relaxed to left invertibility of $\mathbf{P}_{Z\mid U,a}$ and $\mathbf{P}_{W\mid U}$, allowing cases where $Z$ and $W$ each have at least as many levels as $U$. Perhaps surprisingly, Assumption 10 can even be relaxed to allow settings where both proxies $Z$ and $W$ have fewer levels than the latent variable. These generalizations are discussed in Section 3.1.
Techniques for identifying causal targets via array decomposition, presented in works including Kuroki and Pearl (2014), Allman et al. (2015), and Zhou and Tchetgen Tchetgen (2024) can be viewed as special cases of Kruskal’s uniqueness theorem for the Candecomp/Parafac (CP) decomposition (Kruskal, 1977), which we discuss in Section 3.1. As a consequence, any factor of the full law of hidden and observed variables can be recovered in these settings up to latent category labels. A natural extension of these ideas to continuous settings is explored in Section 3.2.
@@ 3.1 Discrete Hidden Variable Models
The array decomposition approach in discrete models is a special case of Kruskal’s uniqueness theorem for the Candecomp/Parafac (CP) decomposition of a three-way array (i.e., third-order tensor), as noted in Allman et al. (2015). The theorem gives general conditions for uniqueness of a three-way array decomposition. Applied to the distribution of the observed variables, it provides conditions under which any factorization of the full law of hidden and observed variables is identified up to latent state labels, and rules out representations with a lower-dimensional latent variable.
The remainder of this section presents the theorem and examines how its assumptions relate to those of Kuroki and Pearl (2014). A concise proof of the theorem is provided by Rhodes (2010), and the theorem is discussed at length in Stegeman and Sidiropoulos (2007) and Derksen (2013).
@@
The rank-$R$ CP decomposition of a three-way array $\boldsymbol{\mathcal{T}}\in\mathbb{R}^{M\times N\times J}$ consists of factor matrices $\mathbf{A}\in\mathbb{R}^{M\times R}$, $\mathbf{B}\in\mathbb{R}^{N\times R}$, and $\mathbf{C}\in\mathbb{R}^{J\times R}$ such that, for each $j=1,\dots,J$, the matrix obtained by fixing the third index of $\boldsymbol{\mathcal{T}}$ at $j$,
$\boldsymbol{\mathcal{T}}_{j}:=\boldsymbol{\mathcal{T}}(:,:,j),$
admits the representation (Stegeman and Sidiropoulos, 2007; Kruskal, 1977)
$\boldsymbol{\mathcal{T}}_{j}=\mathbf{A}\mathbf{C}_{j}\mathbf{B}^{\top},$ (7)
where $\mathbf{C}_{j}$ is a diagonal matrix whose diagonal entries are given by the $j^{\text{th}}$ row of $\mathbf{C}$.
@@
The three-way rank of $\boldsymbol{\mathcal{T}}$ is the smallest integer $R$ for which a CP decomposition of the form in Equation (7) exists.
@@
The Kruskal rank (or $k$-rank) of a matrix $\mathbf{M}$, denoted $k_{M}$, is the largest integer $k$ such that any subset of $k$ columns of $\mathbf{M}$ is linearly independent.
@@
A three-way array $\boldsymbol{\mathcal{T}}$ has three-way rank R, and its rank-$R$ CP decomposition is essentially unique, if:
@@
where $\operatorname{diag}(f_{U\mid a})$ denotes the diagonal matrix with $(i,i)$-th entry $\Pr(U=u_{i}\mid A=a)$. In Definition 3.1, this yields
$\boldsymbol{\mathcal{T}}_{j}=\mathbf{A}\,\mathbf{C}_{j}\,\mathbf{B}^{\top}=\mathbf{P}_{y_{j},W,Z\mid,a}=$
@@
Invoking Theorem 3.5, identification of latent factors
$f_{W\mid U}(w\mid u_{i}),f_{Y\mid U,A}(y\mid u_{i},a),\text{ and }f_{U,Z\mid A}(u_{i},z\mid a)$
together reduces to solving the tensor decomposition separately for each $a\in\mathcal{A}$. The decomposition identifies the latent categories only up to a permutation, so the index $i$ serves merely as a label for the latent state $u_{i}$. Since component $f_{W\mid U}(w\mid u_{i})=f_{W\mid U}(w\mid u_{i},a)$ for all values of $a\in\mathcal{A}$, labels can be aligned across treatment levels. Hence, $f_{A,Y,U,W,Z}(a,y,u_{i},w,z)$ is identified up to a permutation of the latent labels $i$.
The original proof of Theorem 3.5 is nonconstructive (Kruskal, 1977), and can be recast in terms of uniquely determined eigenspaces (Rhodes, 2010). In the special setting considered by Kuroki and Pearl (2014), the authors showed latent components can be obtained directly via eigendecomposition tasks. Modern methods for recovering latent factors under the general assumptions of Theorem 3.5 often rely on iterative procedures such as alternating least squares or gradient-based optimization (Uschmajew, 2012; Kolda and Hong, 2020).
Notably, the inequality in Theorem 3.5 treats $W$, $Z$, and $Y$ symmetrically. Consequently, the condition may still hold even if one observed variable (e.g., $W$) has fewer levels than the latent variable $U$, provided that the other two variables ($Z$ and $Y$) compensate. However, a necessary requirement for the inequality to hold is that $W$, $Z$, and $Y$ together have at least $2|\mathcal{U}|+2$ distinct categories.
The CP decomposition framework is typically introduced in the simpler setting of a single latent variable with three observed variables that are mutually independent conditional on the latent (e.g., Figure 3 omitting $A$). The three-way array formed by the observed variables uniquely determines the full law up to latent state labels under the condition in Theorem 3.5. We now turn to how this framework extends to continuous models.
@@ 3.2 Continuous Hidden Variable Models
When the variables in Figure 3 (except treatment $A$) are continuous, Theorem 1 of Hu and Schennach (2008) provides a natural analogue of the array decomposition framework for uniquely determining the full law up to latent state labels. Originally developed for a setting with only three observed variables, Deaner (2023) extends the Hu and Schennach (2008) result to the causal model in Figure 3, identifying the full law $f_{A,Y,U,W,Z}(a,y,u_{i},w,z)$ and thereby counterfactual distribution $f_{Y(a)}(y)$.
The extension relies on the conditional independences in Assumptions 89, the regularity condition in Assumption 12, and additional conditions Assumptions 13 and 14.
@@
$(W,Z,Y,U,A)$ admit a joint density that is bounded and nonzero.
@@
For each $a\in\mathcal{A}$ and $g$ square-integrable, $\mathbb{E}[g(u)\mid w,a]=0\,\text{ a.s.}\iff g(u)=0\,\text{ a.s.}$ and $\mathbb{E}[g(u)\mid z,a]=0\,\text{ a.s.}\iff g(u)=0\,\text{ a.s.}$
@@
For each $a\in\mathcal{A}$ and $i\neq j$, the set
$\{y:f_{Y\mid U,A}(y\mid u_{i},a)\neq f_{Y\mid U,A}(y\mid u_{j},a)\}$
has positive probability under the marginal of $Y\mid A=a$.
Recalling the relationship between completeness and invertibility discussed in Section 2.1, Assumption 13 is analogous to Assumption 10 in Kuroki and Pearl (2014). Additionally, Assumption 14 is analogous to Assumption 11. Thus, the results of Hu and Schennach (2008) and Deaner (2023) can be viewed as natural extensions of Kuroki and Pearl (2014) to the continuous setting.
3.3 Identifying Latent Labels
In the hidden confounding setting considered in Figure 3, label recovery for the latent confounder $U$ is not required for identification of counterfactual distributions or causal effects. The confounder is not the target of intervention; it is integrated out in the adjustment formula (Equation 1, with $U$ in place of $C$). Any relabeling of the latent categories of $U$ leaves the functional (i.e., $\int f_{Y\mid A,U}(y\mid a,u)\,f_{U}(u)\,du$) unchanged.
In contrast, outside of the hidden confounding context, it is often necessary to impose additional conditions in order to recover the labels of the latent variable. For example, when treatment $A$ itself is unobserved but proxies of it are available, allowing identification of $f_{Y\mid A,C}(y\mid a_{i},c)$ (Zhou and Tchetgen Tchetgen, 2024), recovering which $i$ corresponds to $a_{1}$ versus $a_{0}$ is essential to identify a specific counterfactual distribution
$f_{Y(a_{i})}(y)=\int f_{Y\mid A,C}(y\mid a_{i},c)\,f_{C}(c)\,dc,$
because intervention is defined relative to a particular treatment level.
3.3.1 Conditional Unbiasedness
Hu and Schennach (2008) show that in the model of Figure 3, under Assumptions 89 and 1214, latent state labels can be recovered under an unbiasedness condition: a known functional of an observed variable (e.g., a measure of central tendency) evaluated conditional on each latent value equals the latent value itself (Assumption 15).
@@ 3.3.1 Conditional Unbiasedness
There exists a known functional $M$ such that for each $u\in\mathcal{U}$,
$M[f_{W\mid U}(w\mid u)]=u.$
3.3.2 Conditional Monotonicity
When $U$ is ordinal and its support known, an alternative condition, Assumption 16, suffices. Assumption 16 requires that a known functional of an observed variable is strictly ordered across latent categories; since the full joint law is identified, this ordering determines the labels.
@@ 3.3.2 Conditional Monotonicity
$U$ is ordinal with known support and there exists a known functional $M$ such that, for all $u_{i}<u_{j}$
$M[f_{W\mid U}(w\mid u_{i})]<M[f_{W\mid U}(w\mid u_{j})].$
In both Assumption 15 and 16, $W$ may be replaced by $Z$ or $Y$, provided the condition holds additionally conditional on each $a\in\mathcal{A}$.
@@ 4 Model Comparison
The array decomposition approach in Section 3 can be used to recover counterfactual quantities across a broad class of hidden variable models. For example, Zhou and Tchetgen Tchetgen (2024) apply the method to hidden treatment models. This flexibility underscores a key strength of the array decomposition framework. Moreover, the two proximal strategies for hidden confounding invoke distinct conditional independence restrictions, which are non-nested (Deaner, 2023).
We now compare the two approaches under shared conditional independences compatible with both frameworks. Specifically, we consider the hidden confounding model represented in Figure 3, i.e., reflecting the conditional independences in Assumptions 8 and 9.
Table 1: Comparison for discrete hidden confounding model in Figure 3.
@@
$\mathbf{P}_{Z\mid U,a}$ is left invertible $\exists\mathbf{H}$ s.t. $\mathbf{P}_{Y\mid Z,a}=\mathbf{H}\,\mathbf{P}_{W\mid Z,a}$ $\mathbf{P}_{Z\mid U,a}^{-1}$ and $\mathbf{P}_{W\mid U}^{-1}$ exist Some row of $\mathbf{P}_{Y\mid U,a}$ has distinct entries $\mathbf{P}_{Z\mid U,a}^{-1}$ and $\mathbf{P}_{W\mid U}^{-1}$ exist Some row of $\mathbf{P}_{Y\mid U,a}$ has distinct entries
@@
$\mathbf{P}_{Z\mid U,a}$ is left invertible $\exists\mathbf{H}$ s.t. $\mathbf{P}_{Y\mid Z,a}=\mathbf{H}\,\mathbf{P}_{W\mid Z,a}$ $2|\mathcal{U}|+2\leq k_{A}+k_{B}+k_{C}$ $\mathbf{P}_{Z\mid U,a}$ is left invertible
@@
$\exists\mathbf{H}$ s.t. $\mathbf{P}_{Y\mid Z,a}=\mathbf{H}\,\mathbf{P}_{W\mid Z,a}$ $|\mathcal{U}|+2\leq k_{A}+k_{C}$
@@
Bridge equation approach assumptions are from Miao et al. (2018).
@@
4.1 Discrete Model Comparison
For discrete models, we compare Assumption 6 and 7 from the bridge equation approach in Miao et al. (2018) to two alternative assumption sets (Table 1). Set 1 is from Kuroki and Pearl (2014); Set 2 is from Kruskal (1977), the generalization of Set 1 via Theorem 3.5.
Comparing to Set 1, all model overlap occurs under the assumptions of Kuroki and Pearl (2014). In particular, the invertibility conditions in Assumption 10 from Kuroki and Pearl (2014)—namely, that $\mathbf{P}_{Z\mid U,a}^{-1}$ and $\mathbf{P}_{W\mid U}^{-1}$ exist—also imply the existence of $\mathbf{P}_{W\mid Z,a}^{-1}$. This directly satisfies Assumptions 6 and 7 in Miao et al. (2018), since one can define
$\mathbf{H}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{-1}\quad\text{so that}\quad\mathbf{P}_{Y\mid Z,a}=\mathbf{H}\,\mathbf{P}_{W\mid Z,a}.$
Thus, the model from Kuroki and Pearl (2014) is contained within that of Miao et al. (2018). In other words, Kuroki and Pearl (2014) imposes stronger assumptions.
Comparing to Set 2, again let
$\mathbf{A}=\mathbf{P}_{W\mid U},\quad\mathbf{B}=\mathbf{P}_{Z\mid U,a}\,\operatorname{diag}(f_{U\mid a}),\quad\mathbf{C}=\mathbf{P}_{Y\mid U,a},$